OpenStudy (anonymous):

Evaluate the integral -3sinx(-4x) dx.

zepdrix (zepdrix):

\[\Large\rm \int\limits -3 \sin x (-4x)~dx\]Is it like this? Or is the -4x supposed to be the argument of the sine? I couldn't tell if that was a typo or not.

OpenStudy (anonymous):

yes

zepdrix (zepdrix):

\[\large \sin x (-4x) \qquad or\qquad \sin(-4x)\]